A two pole three phase 50 Hz, Y – conncted AC generator has 9 slots per pole and 6 conductors per slot. What is the flux per pole if the voltage on open circuit is 1.1kV. Assume a coil span of unity. 


1
Expert's answer
2021-07-05T04:54:52-0400

"K_d=\frac{\sin{\frac{m\beta}{2}}}{m\sin{\frac{\beta}{2}}}\\
m=\frac{9}{3}=3\\
\beta=\frac{180\degree}{n}\\
n=9\\
\beta=\frac{180\degree}{9}\\
\beta=20\degree\\
K_d=\frac{\sin{\frac{3\times 20\degree}{2}}}{3\sin{\frac{20\degree}{2}}}=0.96\\
Z=9\times \frac{6}{3}=18\\
K_c=1\\
f=50Hz\\
E=1.1kV=1100V\\
\Phi=\frac{E}{2.22\times K_c\times K_d\times Z\times f\times \sqrt{3}}\\
\Phi=\frac{1100}{2.22\times 1\times 0.96\times 18 \times 50\times \sqrt{3}}=0.33Wb\\"